OpenStudy (anonymous):

Logs! Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ? x = –3 x = 6/5 x = 2 x = 4

OpenStudy (anonymous):

put it as 4^(left side) = 4^(right side) this would get rid of the logs and you should be able to solve it quite easily afterwards.

OpenStudy (anonymous):

wait....did you mean log base 4?

OpenStudy (anonymous):

if not, use 10^(left side) = 10^right side

OpenStudy (anonymous):

Yes, log base 4

OpenStudy (anonymous):

\[f(x) = \log_{4} x + \log_{4} (x-3) = 1\] This

OpenStudy (anonymous):

4^log4 x* 4^log4 (x-3) = 4^1 x (x-3) = 4

OpenStudy (anonymous):

So, x = 4?

OpenStudy (anonymous):

:) yep.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@Puzzhang, Be careful If x (x-3) = 4, then \[x^2-3x-4 =0\] \[(x-4)(x+1)=0\] Then x=4 or x= -1 @rebercca14

OpenStudy (anonymous):

true. but since it was a multiple choice question.....agreed there would be both those answers in any other circumstance though.

OpenStudy (anonymous):

I mean the exact answer :D the x=-1 is not in the options. Just saw now. make sure you don't teach this way because if it is multi choice multi correct option, then the person whom you taught this will never know two solutions exist

OpenStudy (anonymous):

Also x cannot be -1 as log (negative quantity) is not posible. So only x=4 is the solution